# Promptlinecolor=black,linewidth=2pt, leftmargin=1cm,rightmargin=1cm MALLM-GAN: Multi-Agent Large Language Model as Generative Adversarial Network for Synthesizing Tabular Data

### v3

@@ Abstract
In the era of big data, access to abundant data is crucial for driving research forward. However, such data is often inaccessible due to privacy concerns or high costs, particularly in healthcare domain. Generating synthetic (tabular) data can address this, but existing models typically require substantial amounts of data to train effectively, contradicting our objective to solve data scarcity. To address this challenge, we propose a novel framework to generate synthetic tabular data, powered by large language models (LLMs) that emulates the architecture of a Generative Adversarial Network (GAN). By incorporating data generation process as contextual information and utilizing LLM as the optimizer, our approach significantly enhance the quality of synthetic data generation in common scenarios with small sample sizes. Our experimental results on public and private datasets demonstrate that our model outperforms several state-of-art models regarding generating higher quality synthetic data for downstream tasks while keeping privacy of the real data.
@@ 1 Introduction
Recent advancements in generative models, particularly in text and image,(7; 28) have shown the benefits of technology for generating synthetic data that resembles real data. Despite this potential, generating tabular data has not been fully tapped into; it has evolved through traditional statistical approaches, like Bayesian networks (38), to deep learning techniques, including autoencoders and Generative Adversarial Networks (GANs) (42). However, these methods require large amounts of data for their training, contradicting our objective of solving data scarcity. Also, the sparsity and heterogeneity in tabular data make GAN or other deep learning a suboptimal choice, as evidenced by the fact that the tree-based method (e.g., XGBoost) works better than deep learning model (11).
@@
Recently, advancements in large language models (LLMs) have also enabled researchers to use their general intelligence to synthesize tabular data.(6; 14) The premise is that prior knowledge encoded in the parameters of LLMs can provide contextual knowledge for coherent semantics that is required to learn the underlying data generation process. Several studies transformed tabular data to natural language via serialization, and used pre-trained LLMs to generate text containing the synthetic tabular data (6; 14; 21). While fine-tuning LLMs has led to the creation of more nuanced synthetic data, this process again requires a significant sample size, contradicting the objective of addressing data scarcity.
@@
In contrast, in-context learning presents a promising alternative by allowing LLMs to customize without compromising their general reasoning abilities (24) . Particularly, few-shot learning in in-context learning is to provide a few “examples” of data to let LLM to learn the patterns and mimic the examples (17). Our study aims to utilize this few-shot capability for synthetic tabular data generation. Currently, tabular generative model by in-context learning, such as (31) faces a critical shortcoming; it only accommodate too few “examples” (not an entire dataset available), thus discarding remaining data that cannot manage to fit into limited context length (input token). For example, in our real-world trial data ATACH2 with 37 variables, just ten samples consume 4,232 input tokens, whereas the common input token size of LLMs, such as GPT3, is 2,048. This failure to utilize all available scarce data can make LLMs perceive the underlying data-generating process merely based on “educated guess” with its prior knowledge instead of the data itself. This, consequently, causes distribution discrepancy between real data and synthetic data.
@@
Therefore, we aim to bridge this critical gap. Our key idea is to make the data generation process explicit; the objective of our in-context learning is to generate a better data generation process, as well as to generate individual data instances. Here, the data generation process is a prompt text that consists of the context of data and any simple “model” that describes the relationship between data variables. We chose to use a Bayesian network or causal structure due to its interpretability and simplicity.
However, another challenge is to identify the ground-truth data generation process. Motivated by GAN’s adversarial training, we optimize the data generation process (“generator”) in adversarial training with “discriminator” (Table 1). The discriminator’s role is to discriminate real data from the generated data, and we use the accuracy of the discriminator as a “loss function” to be minimized to optimize the generator. Unlike GAN, our generator is a text format, which doesn’t have derivatives. We address it by Optimization by Prompting, which leverages an independent LLM as an optimizer (43). After optimizing the data generation process, the LLM as generator uses it to finally generate synthetic data.
@@
Contribution of this paper can be summarized as below
@@ 2 Related Work
This section provides a brief overview of the most relevant prior work in the field of tabular data generation, LLM-supported synthetic data generation, and the different roles that LLM plays in real-world applications.
Synthetic tabular data generation. Many studies have been proposed to generate high-quality tabular data for privacy-preserving data sharing and augment training data size for machine learning models. Traditional method including Bayesian network (44; 38), approximation Bayesian computation (5), and SMOTE (8). Particularly, a Bayesian network can represent a pairwise causal structure via directed acyclic graph (DAG) (26), in which a directed edge between variable A and B exists if A causes B. Causal structure is a compact representation of underlying variable relationship, but does not fully capture the nonlinear numerical relationship between the different types of variables. Deep generative models have also been widely utilized. This includes variational autoencoders such as TVAE (42) and others (2; 40); GAN such as CTGAN (42), and diffusion model such as TabDDPM (19). The limitation of these models is that they require a sufficient sample size to train the generative model, which contradicts our aim to solve data scarcity.
Synthetic data generation using LLMs. Recently, large language models emerged and have demonstrated their powerful performance in generating natural text. It has also shown a great potential in tabular data (10), such as for predicting (14; 12; 45; 21) and generating tabular data (6; 32; 47; 12). Among them, GReaT (6), the first model in this line, transformed the tabular data into text and fine-tuning LLM (GPT-2), including a feature order permutation step for added realism. However, these prior methods require fine-tuning LLMs, which requires large data size and computing resources for fine-tuning. This limitation of existing models motivates us to develop few-shot tabular data generative model.
Multiple roles of LLM in real-world applications. In addition to typical natural text generation tasks it was originally trained for, LLMs have been utilized in various tasks. LLM has been used as an optimizer for the data type that we can not calculate derivatives directly (43). This LLM as optimizer was used to optimize prompt (43) and optimize heuristic algorithms written in codes (29). LLMs also have been utilized for communicating with other LLMs (i.e., multi-agent LLM); multiple LLMs play different roles to accomplish a task collaboratively, such as coding,(13) question answering,(41) and online decision making.(35; 16)
LLM and causal structure discovery Causal structure discovery involves either data-driven or expert knowledge-driven approaches. Data-driven approaches discover causal structure from data by conditional independence tests (33; 34), score-based heuristics (36), or relaxing the discrete constraints into continuous optimization (48; 46). Despite these advances, identifying the ground-truth causal structure from data remains a significant challenge, particularly in complicated domains like healthcare or when data is scarce. Expert-driven approaches can be an alternative option, but these methods are time-consuming and require significant expert involvement. Recently, many studies (18; 23) suggest that LLMs, which encode prior knowledge in their parameters, can support causal discovery by complementing expert knowledge. In this paper, we leverage multiple LLMs with different roles to mimic adversarial training in GAN and use the heuristic causal structure discovery to guide the data generation process.
@@ 3.2 Multi-agent LLM as GAN
Overview. We propose to develop a multi-agent LLM as GAN (MALLM-GAN) that generates tabular data by mimicing adversarial optimization (Fig. 1). The objective is to optimize the data generation process , which is a natural language description of i) the problem description and ii) the simple data generation process or causal structures representing relationships between variables. In MALLM-GAN, for each iteration , an LLM agent Generator generates data with and a batch in ; a supervised model Discriminator is accordingly optimized using and evaluates using ; and another LLM agent Optimizer improves to decrease the discriminator’s accuracy (Algorithm 1). We repeat the iterations until the discriminator’s accuracy converges or the iteration reaches the maximum epoch.
@@ 3.2.1 Generator
Note that subscription for iteration will be omitted for simplicity without loss of generalizability. Also, note that we used causal structure as a means to convey the relationship between variables within the prompt; thus, obtaining ground-truth causal structure is not our ultimate goal.
Few shot examples. The data generation process is supplemented with examples to leverage in-context few-shot learning. Structured data is serialized into JSON format, e.g., {age: 53, work class: self-emp, …} (Supplement listing 2 Lines 25-28). Various natural language serializations were tested but had minimal impact on performance. The number of examples is crucial; a large allows learning from diverse examples but is constrained by context length, while a small avoids overflow but underutilizes data. Our solution, “batches in a batch,” splits a batch into smaller pieces that fit the input token size, generates a set of synthetic data, and collates them into (see Algorithm 1 Line 6). This approach balances the trade-offs in in-context few-shot learning. The final input to the generator LLM is:
for each optimization iteration. See Supplement listing 2 for a full example.
LLM as generator. With the prompt in Eq. 2, the pre-trained, frozen LLM (e.g., GPT-3.5) generates synthetic data. The goal is to create similar but not identical text to the samples, with the temperature parameter controlling variability. The temperature is set low enough to maintain the original data distribution but high enough to avoid copying. The generator LLM runs multiple times with smaller examples in a batch, and the generated data is collated into . denotes the synthetic data generated at iteration . See Supplement listing 3 for an example.
Conditional sampling. As MALLM-GAN generates synthetic data using LLM, it seamlessly inherits the benefits of LLM, such as conditional sampling. LLM predicts the next tokens given a user-provided context, even when the specified condition is rare or given as a range. To conditionally sample the synthetic data, we modify the task instruction to contain specific conditions (Supplement Listing 4).
@@ 3.2.3 Optimizer
The next task is to optimize based on its score . Our parameter to optimize is , a text, which doesn’t have derivatives. So we use Optimization by Prompting, which leverages LLM as an optimizer (43). To make LLM acts as a optimizer, we provide a meta-prompt, which consists of the causal structure and the optimization task descriptions such as “Your task is to optimize prompts for generating high-quality synthetic data. Aim … ” (see Example in Supplement listing 5 Line 3-6).
To leverage LLM’s in-context few-shot learning in the optimizer (43), we provide a few “examples” of possible solutions (). Note that the example here is different from data . We keep the top solution pairs over the past iteration as the optimization trajectory to guide the optimization. We sort the score, so that the more desirable goes to the end of prompt. This will allow the LLM to recognize patterns among the data generation process with better score. See example in Supplement listing 5 Line 9-31.
A potential pitfall is that the of past iteration is not comparable to the of current iteration . The past discriminators have much lower performance in discriminating real and fake, thus the score are not reliable to compare of past iterations with of current iteration. Thus we adjust the score of past iterations with the latest discriminator , so that all the scores are directly comparable to select best .
In all, the optimizer LLM takes as input the meta prompt and a series of data generation process and adjusted scores . The optimizer outputs the revised data generation process, particularly focusing on causal structure and task instruction. We repeat the iterative optimization and generation until reaching to the maximum iteration.
@@ 4 Experiments
We present the evaluation results of MALLM-GAN. Our extensive experiments demonstrated that MALLM-GAN outperforms baselines in generating high-quality synthetic data while preserving data privacy, thanks to the adversarial optimization of the data generation process. Additionally, MALLM-GAN’ provides explainable data generation through natural textual representation, effectively generating high-quality synthetic data based on user-provided conditions, even for rare categorical values or numeric ranges.
@@ 4.1 Setting
LLM. We used HIPPA-compliant Azure OpenAI GPT-3.5(7) as our generator and GPT-4 (25)(gpt-4-32k-0613) as our optimizer. Due to the extensive workload of the generator, we opted for the lighter, faster gpt-35-turbo-0125 model with a 16k context length. For the optimizer, requiring combinatorial search and high-level reasoning, we used the up-to-date gpt-4-32k-0613 model. As the optimizer requires more “creativity” than the generator, the generator’s temperature was set to 0.5, and the optimizer’s to 1.0 after multiple trials.
Discriminator. Strong discriminators do not always contribute to a better generator (3). We tested Logistic regression, XGBoost, and neural network; we used the logistic regression model because it showed the highest performance while ensuring tractability during incremental updates over the iterations (Supplementary 6).
@@
Baselines. We compared MALLM-GAN with multiple state-of-the-art tabular generative models such as: traditional over-sampling techniques, SMOTE (9), the variational auto-encoder, TVAE (42), the generative adversarial network, CTGAN (42), LLM-based synthetic data generation model, Be-GReaT(6), and a diffusion model, TabDDPM (19). Similar to MALLM-GAN, a prior work (31) uses in-context few-shot learning of pre-trained LLMs but incorporates post-hoc data selection, which is beyond our scope. A comparison without post-hoc selection is available in Table 3.
Other hyperparameters. Various serialization techniques have been proposed to transform tabular data into natural language text (14). We tested several serializations, such as Manual Template (14) (“Age is 53, Work class type is self-employed, …”), which proved ineffective for moderate feature sizes; this serialization made the input prompt lengthy and talkative, only worked when feature size is very small. Feature order permutation (6) also had negligible impact on performance. Specific hyperparameters and computing resources are available in Supplement Section 2.
Training data size vs. quality of synthetic data. We evaluated the impact of training data size on synthetic data quality by sampling subsets of different sizes (). We particularly aimed to compare performances in low and moderate data size. For fair comparison between real and synthetic data, synthetic data was generated to match the size of real data (). We held out 200 samples as the test set before sampling and replicated experiments for each sub sample five times to estimate the standard error of the evaluation metrics. The batch size was set to be 50, with maximum iterations set to be 5, 4, 3, 2 for data sizes , respectively.

### v4

# MALLM-GAN: Multi-Agent Large Language Model as Generative Adversarial Network for Synthesizing Tabular Data

@@ Abstract
In the era of big data, access to abundant data is crucial to driving research forward. However, such data are often inaccessible due to privacy concerns or high costs, particularly in the healthcare domain. Generating synthetic (tabular) data can address this, but existing models typically require substantial amounts of data to train effectively, contradicting our objective of solving data scarcity. To address this challenge, we propose a novel framework to generate synthetic tabular data, powered by large language models (LLMs) that emulate the architecture of a Generative Adversarial Network (GAN). By incorporating the data generation process as contextual information and utilizing LLM as the optimizer, our approach significantly enhances the quality of synthetic data generation in common scenarios with small sample sizes. Our experimental results on public and private datasets demonstrate that our model outperforms several state-of-art models regarding generating higher quality synthetic data for downstream tasks while keeping the privacy of the real data. Code is available at: https://anonymous.4open.science/r/MALLM-GAN-1F5B.
@@ 1 Introduction
Recent advancements in generative models, particularly in text and image,(7; 30) have shown the benefits of technology for generating synthetic data that resembles real data. Despite this potential, tabular data generation has not been fully tapped into; it has evolved through traditional statistical approaches, such as Bayesian networks (38), to deep learning techniques, including autoencoders and Generative Adversarial Networks (GANs) (41). However, these methods require large amounts of data for their training, contradicting our objective of solving data scarcity. Furthermore, the sparsity and heterogeneity in tabular data make GAN or other deep learning a suboptimal choice, as evidenced by the fact that the tree-based method (e.g., XGBoost) works better than the deep learning models(11).
@@
Recently, advancements in large language models (LLMs) have also enabled researchers to use their general intelligence to synthesize tabular data.(6; 14) The premise is that prior knowledge encoded in the parameters of LLMs can provide contextual knowledge for coherent semantics that is required to learn the underlying data generation process. Several studies transformed tabular data to natural language via serialization, and used pre-trained LLMs to generate text containing the synthetic tabular data (6; 14; 24). While fine-tuning LLMs has led to the creation of more nuanced synthetic data, this process again requires a significant sample size, contradicting the objective of addressing data scarcity.
@@
In contrast, in-context learning presents a promising alternative by allowing LLMs to customize without compromising their general reasoning abilities(27). In particular, a few-shot learning in in-context learning is to provide a few “examples” of data to allow LLM to learn the patterns and mimic the examples (19). Our study aims to utilize this few-shot capability for synthetic tabular data generation. Currently, a tabular generative model by in-context learning, such as(33), faces a critical pitfall that it cannot accommodate entire table due to context window size. This failure to utilize all available scarce data can make LLMs perceive the underlying data-generating process merely based on “educated guess” with its prior knowledge instead of the data itself. This, consequently, causes distribution discrepancy between real data and synthetic data.
@@
Therefore, we aim to bridge this critical gap. Our key idea is to make the data generation process explicit; the objective of our in-context learning is to generate a better data generation process, as well as to generate individual data instances. Here, the data generation process is a prompt text that consists of the context of data and any simple “model” that describes the relationship between data variables.
@@
However, another challenge is to identify the ground-truth data generation process. Motivated by GAN’s adversarial training, we optimize the data generation process (“generator”) in adversarial training with “discriminator” (Table 1). The discriminator’s role is to discriminate real data from the generated data, and we use the accuracy of the discriminator as a “loss function” to be minimized to optimize the generator. Unlike GAN, our generator is a text format, which does not have derivatives. We address it by Optimization by Prompting, which leverages an independent LLM as an optimizer (42). After optimizing the data generation process, the LLM as generator uses it to finally generate synthetic data.
The contributions of this paper can be summarized as below:
@@ 2 Related Work
Synthetic tabular data generation. A wide range of methods have been proposed to generate high-quality tabular data for privacy preservation and data augmentation. Traditional approaches include Bayesian networks (43; 38), Approximate Bayesian Computation (4), and SMOTE (8). Deep generative models—such as VAEs (e.g., TVAE (41)) and GANs (e.g., CTGAN (41))—have gained popularity, along with recent diffusion-based models like TabDDPM (22), StaSy (35), and Tabsyn (45). These models, however, typically require large datasets for training, which limits their utility in low-data scenarios. TabPFN (15; 16), a transformer-based model designed for small tabular datasets, shows promise for synthetic generation but struggles with high-cardinality categorical features.
Synthetic data generation using LLMs.Large language models (LLMs) have demonstrated strong capabilities in tabular tasks, including prediction (14; 12; 44; 24) and data generation (6; 34; 46). GReaT (6), for example, fine-tunes GPT-2 on text-transformed tables. However, such approaches require fine-tuning, which is data- and resource-intensive. This motivates few-shot generative methods that avoid full model training. Also, the EPIC framework (21), which leverages LLM to augment imbalanced data for classification, demonstrates LLM’s ability to infer on structured data and perform augmentation.
Multiple roles of LLM in real-world applications. Beyond generation, LLMs have been used to optimize prompts and non-differentiable objectives via heuristic feedback (42; 31). LLMs also operate in multi-agent systems, collaborating to solve complex tasks like coding (13), question answering (40), and decision-making (36; 18). Recent studies suggest that LLMs can also support causal discovery by leveraging encoded domain knowledge (20; 26).
In this paper, we leverage multiple LLMs with different roles to mimic adversarial training in GAN and use the heuristic causal structure discovery to guide the data generation process.
@@ 3.2 Multi-agent LLM as GAN
Overview. We propose to develop a multi-agent LLM as GAN (MALLM-GAN) that generates tabular data by mimicking adversarial optimization (Fig. 1). The objective is to optimize the data generation process , which is a natural language description of i) the problem description and ii) the simple data generation process or causal structures representing relationships between variables.
In MALLM-GAN, for each iteration , an LLM agent Generator generates data with and a batch in ; a supervised model Discriminator is accordingly optimized using and evaluates using ; and another LLM agent Optimizer improves to decrease the discriminator’s accuracy (Algorithm 1). We repeat the iterations until the discriminator’s accuracy converges or the iteration reaches the maximum epoch.
@@ 3.2.1 Generator
Note that subscription for iteration will be omitted for simplicity without loss of generalizability. Also, note that we used the causal structure to convey the relationship between variables within the prompt; thus, obtaining the causal structure of the ground truth is not our ultimate goal.
Few shot examples. The data generation process is supplemented with examples to leverage in-context few-shot learning. Structured data is serialized into JSON format, e.g., {age: 53, work class: self-emp, …} (Supplement listing 1 Lines 25-28). Various natural language serializations were tested but had minimal impact on performance. The number of examples is crucial; a large allows learning from diverse examples, but is restricted by context length, while a small avoids overflow, but underutilizes the data. Our solution, “batches in a batch,” splits a batch into smaller pieces that fit the input token size, generates a set of synthetic data, and collates them into (see Algorithm 1 Line 6). This approach balances the trade-offs in in-context few-shot learning. The final input to the generator LLM is:
for each optimization iteration. See Supplement listing 1 for a full example.
LLM as generator. With the prompt in Eq. 2, the pre-trained, frozen LLM (e.g., GPT-3.5) generates synthetic data. The goal is to create text similar to but not identical to samples, with the temperature parameter controlling the variability. The temperature of the generation is set to 1 to main the original distribution while avoiding hard copying. The generator LLM runs multiple times with smaller examples in a batch, and the generated data is collated into . denotes the synthetic data generated at iteration . See Supplement listing 2 for an example.
Conditional sampling. As MALLM-GAN generates synthetic data using LLM, it seamlessly inherits the benefits of LLM, such as conditional sampling. LLM predicts the next tokens given a user-provided context, even when the specified condition is rare or given as a range. To conditionally sample the synthetic data, we modify the task instruction to contain specific conditions (Supplement Listing 5).
@@ 3.2.3 Optimizer
The next task is to optimize based on its score . Our parameter to optimize is , a text, which doesn’t have derivatives. So we use Optimization by Prompting, which leverages LLM as an optimizer (42). To make LLM acts as a optimizer, we provide a meta-prompt, which consists of the causal structure and the optimization task descriptions such as “Your task is to optimize prompts for generating high-quality synthetic data. Aim … ” (see Example in Supplement listing 3 Line 3-6).
To leverage LLM’s in-context few-shot learning in the optimizer (42), we provide a few “examples” of possible solutions (). Note that the example here is different from data . We keep the top solution pairs over the past iteration as the optimization trajectory to guide the optimization. We sort the score, so that the more desirable goes to the end of prompt. This will allow the LLM to recognize patterns among the data generation process with better score. See example in Supplement listing 3 Line 9-31.
A potential pitfall is that the of the previous iteration is not comparable to the of the current iteration . The past discriminators have much lower performance in discriminating real and fake samples, thus the score are not reliable to compare of past iterations with of current iteration. Thus, we adjust the score of past iterations with the latest discriminator , so that all scores are directly comparable to select best .
In total, the LLM optimizer takes as input the meta prompt and a series of data generation process and adjusted scores . The optimizer outputs the revised data generation process, particularly focusing on causal structure and task instruction. We repeat iterative optimization and generation until we reach the maximum iteration.
@@ 4 Experiments
We present the evaluation results of MALLM-GAN. Our extensive experiments demonstrated that MALLM-GAN outperforms baselines in generating high-quality synthetic data while preserving data privacy, thanks to the adversarial optimization of the data generation process. Additionally, MALLM-GAN provides explainable data generation through natural textual representation, effectively generating high-quality synthetic data based on user-provided conditions, even for rare categorical values or numeric ranges.
@@ 4.1 Setting
LLM. We used HIPPA-compliant Azure OpenAI GPT-4o(28) as our generator and optimizer. Generator’s temperaure of generator was set to be 0.5 to generate data points of highest confidence without randomly guess, while optimizer was set to be more creative using temperature 1.
Discriminator. Strong discriminators do not always contribute to a better generator (2). We tested Logistic regression, XGBoost, and neural network; we used the logistic regression model because it showed the highest performance while ensuring tractability during incremental updates over the iterations (Table 7).
@@
Baselines. We compared MALLM-GAN with multiple state-of-the-art tabular generative models such as: traditional over-sampling techniques, SMOTE (9), the variational auto-encoder, TVAE (41), the generative adversarial network, CTGAN (41), LLM-based synthetic data generation model, Be-GReaT(6), and diffusion models, TabDDPM (22) and Tabsyn (45). We also adapted a transformer based tabular foundation model, TabPFN,(16) to do generation task. Similar to MALLM-GAN, a prior work (33) uses in-context few-shot learning of pre-trained LLMs but incorporates post-hoc data selection, which is beyond our scope. A comparison without post-hoc selection is available in Table 3.
Other hyper-parameters. Various serialization techniques have been proposed to transform tabular data into natural language text (14). We tested several serializations, such as Manual Template (14) (“Age is 53, Work class type is self-employed, …”), which proved ineffective for moderate feature sizes; this serialization made the input prompt lengthy and talkative, only worked when feature size is very small. Feature order permutation (6) also had a negligible impact on performance. Specific hyper-parameters and computing resources are available in Supplement Section B.
Training data size v.s. quality of synthetic data. We evaluated the impact of training data size on synthetic data quality by sampling subsets of different sizes (). We particularly aimed to compare performances in low and moderate data size. For fair comparison between real and synthetic data, synthetic data was generated to match the size of real data (). We held out 200 samples as the test set, and replicated experiments for each subsample five times to calculate the standard error of the evaluation metrics.
5 Results
| Public dataset | Private dataset | ||||||
| Adult () | Magic() | Asia () | Insurance() | ATACH() | ERICH() | ||
| N=25 | Real data | 0.80 | 0.79 | 0.83 | 0.52 | 0.25 | -0.23 |
| SMOTE* | |||||||
| CTGAN | |||||||
| TVAE | |||||||
| Be-GReaT | - | ||||||
| TabDDPM | - | ||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | |||||
| MALLM-GAN | |||||||
| N=50 | Real data | 0.76 | 0.79 | 0.82 | 0.78 | 0.17 | -0.01 |
| SMOTE* | |||||||
| CTGAN | |||||||
| TVAE | |||||||
| Be-GReaT | - | ||||||
| TabDDPM | - | ||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | |||||
| MALLM-GAN | |||||||
| N=100 | Real data | 0.86 | 0.83 | 0.83 | 0.82 | 0.26 | -0.04 |
| SMOTE* | |||||||
| CTGAN | |||||||
| TVAE | |||||||
| Be-GReaT | |||||||
| TabDDPM | - | ||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | - | ||||
| MALLM-GAN | |||||||
| N=200 | Real data | 0.85 | 0.81 | 0.83 | 0.83 | 0.27 | 0.16 |
| SMOTE* | |||||||
| CTGAN | |||||||
| TVAE | |||||||
| BeGReaT | |||||||
| TabDDPM | - | ||||||
| Tabsyn | - | - | |||||
| TabPFN | - | - | - | ||||
| MALLM-GAN | |||||||